# Diffusion and instabilities in large-N holographic Fermi liquids: the vector fluctuations of the electron star

### v2

@@ 5.1 Analytical Green’s functions and instability
which splits the two regimes: [^14] Here we recognize the issue we encountered when we were expanding the equations of motion in small $\epsilon$ in Eqs. (62-67) when we had to deal with the ambiguity of $\gamma$, assuming $\omega\sim|k|^{\gamma}$.
@@
The presence of the linear term might explain why Kaplis:2013gux finds an unexpected temperature dependence of the diffusion constant for certain parameter regimes: in Kaplis:2013gux it is assumed that the dispersion relation describes conventional hydrodynamics, with purely quadratic diffusion. We find instead even more unexpected behavior, with a linear term that can give rise to an instability. On the other hand, Kaplis:2013gux studies also the quasinormal modes, which we do not examine in this work.
@@ 6 Discussion and conclusions
The big mystery is the presence of a pure imaginary pole in the upper frequency half-plane at intermediate temperatures. This signals an instability. Moreover, it is an IR instability (starting at arbitrarily small $k$), which implies that the electron cloud is not a true ground state, and hence likely not a viable holographic dual of a Fermi liquid at finite temperature. Understanding this instability is the main goal for future work in this direction. It might come (1) from the pathologies of the fluid limit – once the non-classical tails and WKB corrections Medvedyeva:2013rpa are taken into account, the system might become stable (2) simply from the existence of some other, thermodynamically preferred, solution of the (nonlinear) Einstein-Maxwell system (12-14). [^19] One possibility is that the true solution is more compact while the electron star has a Jeans instability. This would nicely match with the IR nature of the instability. We thank Sean Hartnoll for this idea. So far we have not been able to confirm either of the above scenarios; we also cannot discard other possibilities. In dual field theory this instability, present for a range of temperatures, could imply that the system corresponds to a quantum critical phase. One could expect that the instability should also manifest itself as a critical point or region in free energy. However, in order to see it one would need to add another operator which presumably condenses (acquires a VEV) at the transition point. Since we do not know what this operator is, we cannot perform this kind of analysis. In other words, we do not know what to compare with the star/cloud free energy. A possible alternative to thermodynamic analysis would be the study of quantum information measures such as those in DiNunno:2021eyf .
@@
An encouraging aspect is the fact that low-$\omega$, low-$k$ expansion can be performed analytically and confirms all the key findings of the numerics, including the instability. This, together with the fact that we have reproduced the known ES-RN phase diagram and the known RN response from Edalati:2010hk ; Davison:2011uk ; Davison:2013bxa , makes us confident in our findings. Achieving a full understanding of the mysterious instability is now an important task for future work. However, it also prompts us to look for a substantially different approach toward holographic (non-)Fermi liquids.
@@ Acknowledgements.
We are grateful to Blaise Gouteraux, Sašo Grozdanov, Hareram Swain, Juan Pedraza and Sean Hartnoll for stimulating discussions. The work on Sections 1, 2 and 6 (backgrounds, supervision and interpretation of results) was supported by Russian Science Foundation Grant No. 24-72-10061 [https://rscf.ru/project/24-72-10061/] and performed at Steklov Mathematical Institute of Russian Academy of Sciences (Mihailo Čubrović). The rest of the paper was funded by the Institute of Physics Belgrade, through the grant by the Ministry of Science, Technological Development, and Innovation of the Republic of Serbia.

### v3

@@ 5.1 Analytical Green’s functions and instability
which splits the two regimes: [^13] Here we again recognize the issue we encountered when we were expanding the equations of motion in small $\epsilon$ in Eqs. (62-67) when we had to deal with the ambiguity of $\gamma$, assuming $\omega\sim|k|^{\gamma}$.
@@
The presence of the linear term might explain why Kaplis:2013gux finds an unexpected temperature dependence of the diffusion constant for certain parameter regimes: in Kaplis:2013gux it is assumed that the dispersion relation describes conventional hydrodynamics, with purely quadratic diffusion. We find instead even more unexpected behavior, with a linear term that can give rise to an instability. On the other hand, Kaplis:2013gux also studies the quasinormal modes, which we do not examine in this work.
@@ 6 Discussion and conclusions
The big mystery is the presence of a pure imaginary pole in the upper frequency half-plane at intermediate temperatures. This signals an instability. Moreover, it is an IR instability (starting at arbitrarily small $k$), which implies that the electron cloud is not the true ground state, and hence likely not a viable holographic dual of a Fermi liquid at finite temperature. Understanding this instability is the main goal for future work in this direction. It might come (1) from the pathologies of the fluid limit – once the non-classical tails and WKB corrections Medvedyeva:2013rpa are taken into account, the system might become stable (2) simply from the existence of some other, thermodynamically preferred, solution of the (nonlinear) Einstein-Maxwell system (12-14). [^18] One possibility is that the true solution is more compact while the electron star has a Jeans instability. This would nicely match with the IR nature of the instability. We thank Sean Hartnoll for this idea. So far we have not been able to confirm either of the above scenarios; we also cannot discard other possibilities. In dual field theory this instability, present for a range of temperatures, could imply that the system corresponds to a quantum critical phase. One could expect that the instability should also manifest itself as a critical point or region in free energy. However, in order to see it one would need to add another operator which presumably condenses (acquires a VEV) at the transition point. Since we do not know what this operator is, we cannot perform this kind of analysis. In other words, we do not know what to compare with the star/cloud free energy. A possible alternative to thermodynamic analysis would be the study of quantum information measures such as those in DiNunno:2021eyf .
@@
An encouraging aspect is the fact that $\omega, low$-$k$ expansion can be performed analytically and confirms all the key findings of the numerics, including the instability. This, together with the fact that we have reproduced the known ES-RN phase diagram and the known RN response from Edalati:2010hk ; Davison:2011uk ; Davison:2013bxa , makes us confident in our findings. Achieving a full understanding of the mysterious instability is now an important task for future work. However, it also prompts us to look for a substantially different approach toward holographic (non-)Fermi liquids.
@@ Acknowledgements.
We are grateful to Blaise Gouteraux, Sašo Grozdanov, Hareram Swain, Juan Pedraza, Sean Hartnoll and Ulf Gran for stimulating discussions. The work on Sections 1, 2 and 6 (backgrounds, supervision and interpretation of results) was supported by Russian Science Foundation Grant No. 24-72-10061 [https://rscf.ru/project/24-72-10061/] and performed at Steklov Mathematical Institute of Russian Academy of Sciences (Mihailo Čubrović). The rest of the paper was funded by the Institute of Physics Belgrade, through the grant by the Ministry of Science, Technological Development, and Innovation of the Republic of Serbia.